OpenStudy (anonymous):

A square pyramid is shown below: A square pyramid is shown. The sides of the square base are labeled 0.7 foot. The height of one of the triangular sides is labeled 8 feet. What is the surface area of the pyramid? 11.69 square feet 3.29 square feet 7.49 square feet 2.24 square feet

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@texaschic101 @Hey1122

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

i got the same thing the formaula is A=a2+2aa2 4+h2

OpenStudy (welshfella):

area of base = 0.7^2 area of 4 sides = 4 * 0.5 * 0.7 * 8